OpenStudy (anonymous):

What is the rate of change and initial value for the linear relation that includes the points shown in the table? x y 1 10 2 8 3 6 4 4 (4 points) Initial value: 12, rate of change: –2 Initial value: 8, rate of change: 2 Initial value: 12, rate of change: 2 Initial value 8, rate of change: –2

Mehek (mehek14):

find the slope \(\dfrac{8-10}{2-1}=\dfrac{-2}{1}=-2\) so -2 is the rate of change

Mehek (mehek14):

plug that into the equation we have \(y=-2x+b\\ 10=-2*1 +b\\ 10 = -2 + b \\ 12 = b\) 12 is the initial value

OpenStudy (anonymous):

thanks

Mehek (mehek14):

yw

OpenStudy (anonymous):

can you help me with anoter

OpenStudy (anonymous):

another

OpenStudy (anonymous):

another

OpenStudy (anonymous):

Identify the initial value and rate of change for the graph shown. (4 points) Initial value: 5.5, rate of change: Initial value: 4, rate of change: Initial value: , rate of change: 4 Initial value: , rate of change: 5.5